OpenStudy (anonymous):

Simplify the expression using the properties of exponents: (3a/b-1)-2

OpenStudy (unklerhaukus):

is that \[({3a\over b-1})^{-2}\]?

OpenStudy (anonymous):

Yes. Sorry, I don't know how to enter exponents in here.

OpenStudy (unklerhaukus):

i would wright x to the power y as x^y

OpenStudy (anonymous):

Ok. Thank you.

OpenStudy (unklerhaukus):

\[({3a \over b-1})^{-2}=({b-1 \over 3a})^2={(b-1)^2 \over (3a)^2}\] \[={b^2-2b+1\over 9a^2}\]

OpenStudy (unklerhaukus):

perhaps this is more simple \[{(b-1)^{2} \over 9a^2}\]

OpenStudy (anonymous):

The computer doesn't like that answer. Ummm...ok.